the width of a rectangle is 6 centimeters less than the lenght. the perimeter is 64 centimeters. find the width.
Answer by edjones(7311)

L=w+6
2(L+w)=P
2((w+6)+w)=64
2(2w+6)=64
4w+12=64
4w=52
w=13 cm
L=19 cm
.
Ed